# IMAgent: Training Multi-Image Vision Agents via End2End Reinforcement Learning

### v2

@@ Abstract
Recent VLM-based agents aim to replicate OpenAI O3’s “thinking with images” via tool use, but most open-source methods limit input to a single image, falling short on real-world multi-image QA tasks. To address this, we propose IMAgent, an open-source vision agent trained via end-to-end reinforcement learning dedicated for complex multi-image tasks. By leveraging a multi-agent system, we generate challenging and visually-rich multi-image QA pairs to fully activate the tool-use potential of the base VLM. Through manual verification, we obtain MIFG-QA, comprising 10k samples for training and evaluation. With deeper reasoning steps, VLMs may increasingly ignore visual inputs. We therefore develop two specialized tools for visual reflection and confirmation, allowing the model to proactively reallocate its attention to image content during inference. Benefiting from our well-designed action-trajectory two-level mask strategy, IMAgent achieves stable tool use behavior via pure RL training without requiring costly supervised fine-tuning data. Extensive experiments demonstrate that IMAgent maintains strong performance on existing single-image benchmarks while achieving substantial improvements on our proposed multi-image dataset, with our analysis providing actionable insights for the research community. Codes and data will be released soon.
@@ 1 Introduction
Figure 1: We compared the attention proportions of newly generated tokens by Qwen2.5-VL-7B (a, b) and IMAgent  (c, d) to the input text, image tokens, and previously generated tokens as inference progresses, under single-image input (a, c) and multi-image input (b, d). Qwen2.5-VL-7B’s attention to image tokens gradually decreases and approaches zero under multi-image input (b); in contrast, IMAgent  (c, d), aided by tool use, maintains dynamic visual attention throughout, even with a reasoning chain twice as long.
The agentic reinforcement learning (RL) paradigm [jiang2024mmsearch, jin2025search, singh2025agentic] enables end2end training of LLM-based agents, transforming language model services from passive response generators into proactive decision-makers [acharya2025agentic, zhang2025landscape]. In the vision domain, OpenAI’s O3 [OpenAIo3] demonstrates how interleaved chains of textual and visual reasoning can transcend the limitations of pure VLMs that merely “see” images, enabling true “thinking with images”—a novel inference behavior that significantly expands model capabilities.
Motivated by this breakthrough, many open-source efforts develop VLM-based agents (vision agents), integrating external visual processing tools with the multimodal understanding of base VLM [fu2025refocus, zhang2025chain, zheng2025deepeyes, wu2025vtool, su2025pixel]. By interleaving tool outputs with generated text tokens, these systems partially replicate O3’s advanced reasoning patterns and improve performance on complex tasks.
However, current vision agents are primarily trained to use tools in single-image scenarios [zheng2025deepeyes, wu2025vtool, cao2025ground, zhang2025chain, su2025openthinkimg]. Multi-image tasks require explicit identification and indexing of target images for each operation; thus, directly transferring such models to multi-image settings often fails to meet real-world demands where multi-image understanding is common and essential [liu2024mibench]. On the other hand, existing work focuses mainly on stabilizing tool use, either by carefully curating trajectory data for supervised fine-tuning (SFT) [wang2025simple], modifying RL algorithms [cao2025ground, zheng2025deepeyes, wu2025vtool], or combining both approaches [su2025openthinkimg, su2025pixel, zhang2025chain]. Few studies investigate the underlying principles through which tool use actually enhances agent performance.
To address these issues, we develop IMAgent, the first vision agent trained end2end via agentic RL that can dynamically explore multi-image inputs and solve highly complex visual tasks.
Recognizing that existing multi-image datasets do not sufficiently elicit autonomous tool use behaviors from models, we design a collaborative multi-agent pipeline for data curation. We curate visually rich samples from both natural image [li2025migician] and scientific chart [saxena2025postersum] datasets; generate diverse questions; conduct rigorous manual filtering; and ultimately construct approximately 10k fine-grained QA pairs (9k for training and 636 for evaluation).
Empirical analysis reveals that, without active tool use during chain-of-thought(CoT) reasoning, the model’s attention to input images gradually diminishes, especially in multi-image scenarios. As shown in Figure 1, token-level focus on image content drops to near zero in later inference stages. To mitigate this issue, we introduce two specialized tools—a zoom_in cropping module (visual confirmation) and an image lookback_reuse module (visual reflection), which enable IMAgent to autonomously reallocate attention across relevant regions throughout inference while fostering human-like iterative visual thinking processes.
To avoid costly manual construction of high-quality SFT data while ensuring robust learning dynamics in complex environments, we adopt pure end2end RL training augmented by an action-trajectory two-level mask strategy designed to prevent ineffective trajectories or environment feedback from hindering exploration into advanced forms of visual reasoning. In addition, extensive experiments systematically analyze factors such as tool type that affect the behavior of vision agents during training.
@@
  • We propose IMAgent, the first vision agent aimed at complex single and multi image tasks via end2end RL with MIFG-QA, a multi-image fine-grained dataset we create.
  • To enable dynamic adjustment of image attention throughout reasoning, we design visual reflection and confirmation tools for IMAgent.
  • Extensive experiments show that IMAgent stably uses tools and thinks with images, achieving SOTA results on single and multi-image benchmarks while providing practical training guidelines for the community.
@@ 2.1 Multi-modal Reasoning Models
Multimodal reasoning models show a promising trend toward unifying multiple modalities within a language-centric framework[li2025perceptionreasonthinkplan]. Early works[xu2025llavacotletvisionlanguage, xu2025redstardoesscalinglongcot, yao2024mulberryempoweringmllmo1like] focus on enhancing CoT reasoning through SFT and planning algorithms like Monte Carlo Tree Search[xie2024montecarlotreesearch], improving interpretability and reliability. Building on these foundations, RL-based methods shift the paradigm from imitating static reasoning patterns to actively optimizing them via feedback. Pioneering methods[yu2024rlhfvtrustworthymllmsbehavior, zhang2024improvevisionlanguagemodel, dong2025insightvexploringlongchainvisual] apply Direct Preference Optimization[rafailov2024directpreferenceoptimizationlanguage] for behavioral alignment. After the success of Deepseek-R1[deepseekai2025deepseekr1incentivizingreasoningcapability], Group Relative Policy Optimization (GRPO)[shao2024deepseekmathpushinglimitsmathematical] becomes a mainstream algorithm, and Reinforcement Learning with Verifiable Rewards (RLVR) emerges as a powerful new training paradigm. These methods rapidly expand from mathematical and geometric reasoning[meng2025mmeurekaexploringfrontiersmultimodal, zhan2025visionr1evolvinghumanfreealignment, zhou2025r1zerosahamomentvisual] to visual skills[liu2025visualrftvisualreinforcementfinetuning, liu2025segzeroreasoningchainguidedsegmentation, wang2025vicritverifiablereinforcementlearning], and further to modalities such as video[feng2025videor1reinforcingvideoreasoning] and audio[zhao2025r1omniexplainableomnimultimodalemotion]. Their success highlights the importance of high-quality reward signals, making data augmentation[wu2025synthrlscalingvisualreasoning, liang2025modomodomultidomaindatamixtures], reward design[zhang2025r1rewardtrainingmultimodalreward, wang2025unifiedmultimodalchainofthoughtreward, xu2025mixedr1unifiedrewardperspective, zhang2025viperempoweringselfevolutionvisual] and other modifications[yue2025promotingefficientreasoningverifiable, guo2025ssl4rlrevisitingselfsupervisedlearning] central to performance. Ultimately, RLVR drives breakthroughs in reasoning depth, coherence, and domain adaptability, laying a foundation for training advanced vision agents.
Figure 2: Figure 2: Overview of IMAgent. Our model will automatically choose whether and how to use tools based on the actual problem. Visual verification is achieved by cropping local regions of the image using zoom_in, and visual reflection is achieved by reusing the input image using lookbask_reuse.
@@ 2.2 Vision Agents
Vision Agents aim to break the limitations of static, parameterized knowledge by autonomously using tools, deeply integrating dynamic visual perception with action-based reasoning[chai2025rlfactoryplugandplayreinforcementlearning]. Early approaches relied on prompt engineering to passively trigger tools[jiang2024mmsearchbenchmarkingpotentiallarge, wang2025mllmtoolmultimodallargelanguage, gao2025multimodalagenttuningbuilding, yue-etal-2025-uiorchestra], while current mainstream methods increasingly adopt RL post-training, driving advances in multimodal retrieval[wu2025mmsearchr1incentivizinglmmssearch, geng2025webwatcherbreakingnewfrontier, wang2025vragrlempowervisionperceptionbasedrag] and integrated code execution[gou2024toratoolintegratedreasoningagent, wang2023mathcoderseamlesscodeintegration, li2025torlscalingtoolintegratedrl, feng2025retoolreinforcementlearningstrategic]. Notably, OpenAI’s O3 and Qwen3-VL[QwenTeam2025QwenVL] enable models to “think with images”, generating intertwined chains of thought that combine text and visuals. Inspired by O3’s great performance, many open-source solutions have emerged[zheng2025deepeyes, cao2025ground, zhang2025chain, lai2025minio3scalingreasoningpatterns, su2025pixel, zhang2025chainoffocusadaptivevisualsearch], typically adopting a two-stage post-training process: SFT for tool call format learning, followed by RL to enhance visual tool use for complex tasks. However, this approach requires carefully curated SFT datasets and limits generalization. Deepeyes[zheng2025deepeyes] cultivates tool-calling ability through pure reinforcement learning, but is limited to single image input and a single tool type. Though, aforementioned works achieve partial power of O3, they mainly focus on reproducing the tool use ability in single-image scenarios, largely neglecting the more common multi-image input settings. Addressing this gap is one of the main motivations of our work.
@@ 3.1 Reasoning Pattern
Adhering to a dynamic “think–act–iterate” paradigm, IMAgent outputs the final answer after obtaining sufficient information. Specifically, the workflow proceeds as follows:
@@
    Initial Reasoning. Given an image set $I$ and a corresponding question $Q$, IMAgent performs an initial reasoning step. During this stage, the model autonomously decides whether to invoke visual confirmation or reflection. If the model calls a tool, it will output structured tool use tags.
@@
    Visual Functions. When a visual function is triggered, the model emits the function type and its arguments. For visual confirmation via local image cropping, the model outputs an image index to specify the target image, a bounding box to indicate the region to zoom in on, and a label to emphasize the region of interest. For visual reflection via image reuse, the model outputs an image index and a reflection reason, which reinforces the purpose of reflection and enhances logical interpretability.
@@
    Process Reasoning. The returned results are concatenated with the historical trajectory in an appropriate format, forming the input for the current step. The model continues reasoning based on this augmented trajectory.
@@
3.2 Curation of MIFG-QA
To train the model’s dynamic visual reasoning capabilities for multi-image tasks, we design a three-stage data synthesis pipeline (as shown in Figure 3).
Stage 1: Images Selection. Considering the necessity of activating visual capabilities, we exclude low-resolution images and selected data with rich visual content. Specifically, for natural images, we randomly sample high-resolution data from MGrounding [li2025migician], covering image groups with relationships such as difference, contrast, and time. For scientific figures, we constructe the dataset from high-resolution conference posters [saxena2025postersum]. We perform adaptive structural segmentation on individual conference posters, treating different segmented regions as multi-image data. The resulting multi-image dataset maintains strong contextual relevance.
@@ 3.3 End2End Reinforcement Post-training
Table 1: Results of different models on various benchmarks.
Model E2E Param V* HR-Bench MME MIFG-QA
direct relative Avg 4K 8K Avg nature poster Avg
GPT-4o [hurst2024gpt4o] 66.0 59.0 59.0 59.0 45.2 49.26 47.55 48.11
SEAL [wu2024vstar] 7B 75.4
Dyfo [li2025dyfo] 7B 81.2
Zoomeye [shen2025zoomeye] 7B 93.9 85.5 90.6 69.6 69.3 69.45
LLaVA-OneVision [li2024llava] 7B 75.7 75.0 75.4 63.0 59.8 61.4 57.4 37.93 39.49 38.99
Qwen2.5-VL [bai2025qwen2] 7B 79.13 73.68 76.96 70.62 66.75 68.69 57.3 36.94 44.34 41.98
Qwen2.5-VL-CoT [bai2025qwen2] 7B 73.91 76.32 74.86 68.5 62.25 65.38 59.63 36.35 40.65 39.31
Chain-of-Focus [zhang2025chain] 7B 88.0
Deepeyes [zheng2025deepeyes] 7B 90.43 84.21 87.96 75.0 69.75 72.38 64.0 38.75 43.84 38.36
Pixel Reasoner [su2025pixel] 7B 91.3 77.63 85.86 68.75 63.25 66.0 64.41 41.87 45.72 44.49
IMAgent 7B 88.70 88.16 88.48 73.5 71.5 72.5 64.83 50.24 48.73 49.21
@@ 3.3.1 Reward Shaping
Figure 4: Action-Level and Trajectory-Level Mask.
@@ 3.3.2 Two-Level Mask Strategy
Action-Level Mask. After the model issues a tool call request, the execution results of the tool (including the execution status and returned images) are concatenated into the previous context and used as input for the next round of reasoning. However, since this information often deviates from the distribution of existing knowledge, it can lead to abnormal directions in parameter optimization. To mitigate this issue, we apply a mask to information other than the model’s output to prevent it from participating in policy updates.
Trajectory-Level Mask. xue2025simpletir propose that the instability in multi-turn tool-integrated reasoning training within pure text CoT stems from the emergence and accumulation of low-probability tokens, causing distribution shifts in subsequent model outputs and generating invalid trajectories. We observe a similar phenomenon in multimodal CoT and therefore design trajectory-level mask to filter invalid trajectories.
Specifically, we evaluate $N$ trajectories generated by one sample in the GRPO algorithm; if a trajectory is deemed invalid, the entire trajectory is masked to exclude it from policy updates. This approach incurs almost no additional resource consumption as it directly masks generated trajectories instead of generating new valid ones. In this work, trajectories exceeding the maximum number of iterations, exceeding the maximum response length (potentially caused by repetitive model outputs), or producing no answers are classified as invalid. This mask also avoids reward-induced suppression of the model’s long-range exploration.
@@ 4.1 Experimental Setup
Table 2: Ablation experiments: “w/o MIFG-QA” refers to adopting only single-image training data, “w/o lookback” denotes not using visual reflection tools, and “w/o Traj mask” signifies not masking invalid tracks.
Components V* HR-Bench MME MIFG-QA
direct relative Avg 4K 8K Avg nature poster Avg
baseline 79.13 73.68 76.96 70.62 66.75 68.69 57.3 36.94 44.34 41.98
w/o MIFG-QA 87.83 80.26 84.81 72.75 70.5 71.63 64.02 39.90 40.18 40.09
w/o lookback 86.07 76.32 82.20 74.0 71.25 72.63 62.24 48.25 47.98 48.07
w/o Traj mask 86.09 81.59 84.29 73.5 69.75 71.63 64.03 46.31 46.18 46.23
IMAgent 88.70 88.16 88.48 73.5 71.5 72.5 64.83 50.24 48.73 49.21
@@
Hyperparameters. We adopt an end2end reinforcement learning algorithm, GRPO, with Qwen2.5-VL-7B-Instruct as the base model. The average number of pixels per sample in our MIFG-QA (poster) exceeds 9 million, which occupies over 12k context tokens. We therefore set the maximum pixel budget to 4 million. This preserves context space to accommodate more interaction turns. The model is trained on H20 GPUs with a batch size of 256, 8 rollouts per prompt, and a maximum limit of 5 tool interactions. The KL coefficient is set to 0.0, with the maximum input length restricted to 10,480 tokens and the maximum response length to 20,480 tokens. Our single-image training data is sourced from Deepeyes, which includes fine-grained data of natural images and scientific figures, as well as reasoning data. To enable integrated reasoning over multi-image inputs, we incorporate the constructed MIFG-QA dataset. Main experimental results are presented in Table 1, where the values are the average of three independent experiments conducted with the temperature set to 0, and the maximum pixel limit of input images is 16384×28×28.
@@
Baselines and Benchmarks. We conduct a systematic comparison of IMAgent against advanced proprietary models (e.g., GPT-4o [hurst2024gpt4o]), state-of-the-art open-source models (e.g., LLaVAOneVision-7B [li2024llava], Qwen2.5-VL [bai2025qwen2]), and models implementing multimodal CoT (e.g., Deepeyes [zheng2025deepeyes], Pixels Reasoner [su2025pixel], Chain-of-Focus [zhang2025chain]).
@@
For evaluation benchmarks, we perform comprehensive comparisons on three classic single-image fine-grained benchmarks (V* Bench [wu2024vstar], HR-Bench [wang2025hrbench], and MME-RealWorld [zhang2024mme]) and our proposed MIFG-QA -val to validate the effectiveness of IMAgent’s dynamic visual reasoning. The test data of V* Bench and HR-Bench includes ultra-high-resolution images with resolutions ranging from 2k to 8k, where the target objects relevant to the test questions account for a very small proportion of pixels. Such extensive redundant visual information increases the difficulty for VLMs to accurately locate targets and answer questions. MME-RealWorld is a large-scale high-resolution dataset of real-world scenarios, covering 43 scenarios, which focuses on evaluating the perception and reasoning capabilities of multimodal large models.

### v3

@@ Abstract
Recent VLM-based agents aim to replicate OpenAI O3’s ”thinking with images” via tool use, yet most open-source methods restrict inputs to a single image, limiting their applicability to real-world multi-image QA tasks. To address this gap, we propose IMAgent, an open-source visual agent trained with end-to-end reinforcement learning for fine-grained single/multi-image reasoning. During inference, VLMs tend to gradually neglect visual inputs; to mitigate this issue, we design two dedicated tools for visual reflection and confirmation, enabling the model to actively refocus attention on image content. Beyond that, we, for the first time, reveal how tool usage enhances agent performance from an attention perspective. Equipped with a carefully designed two-layer motion trajectory masking strategy and tool-use reward gain, IMAgent acquires an effective tool-use paradigm through pure reinforcement learning, eliminating the need for costly supervised fine-tuning data. To further unleash the inherent tool-usage potential of the base VLM and fill data gaps, we construct a challenging, visually enriched multi-image QA dataset via multi-agent system. Extensive experiments validate that IMAgent achieves SOTA performance across mainstream single and multi-image benchmarks, and our in-depth analysis offers actionable insights for the community. Code and data will be released soon.
@@ 1 Introduction
Figure 1: IMAgent flexibly select target image from multiple inputs for confirmation and reflection.
@@
Figure 2: We compared the attention proportions by Qwen2.5-VL-7B (a, b) and IMAgent  (c, d) to the input text, image tokens, and previously generated tokens as inference progresses, under single-image input (a, c) and multi-image input (b, d). Qwen2.5-VL-7B’s attention to image tokens gradually decreases and approaches zero under multi-image input (b); in contrast, IMAgent  (c, d), aided by tool use, maintains dynamic visual attention throughout, even with a reasoning chain twice as long.
The success of agentic reinforcement learning (RL) [22, 23, 38] is driving vision-language models (VLMs) beyond passive perception and giving rise to Vision Agents [1, 64]. OpenAI’s O3 [34] demonstrates that agents can autonomously invoke visual tools to achieve “thinking with images.” Inspired by this, the open-source community trains tool-augmented agents (e.g., using zoom_in to extract local details) that already achieve strong progress in single-image fine-grained scenarios [13, 68, 74, 51, 39].
However, real-world applications widely require fine-grained understanding across multiple images [59, 30], where models must capture subtle local cues and perform cross-image comparisons to reach correct conclusions. Due to architectural and cognitive mismatches, naively transferring single-image agents to multi-image tasks often causes systematic failures.
At the operational level, existing agents typically assume a Monolithic Visual Context and lack Explicit Visual Indexing [74, 51, 4, 68, 40], making it difficult to specify which visual entity should be manipulated in multi-image settings. At the cognitive level, long-chain reasoning naturally causes models to forget the initial visual input as text generation grows. Intensive inspection of local details in one image further consumes the context window, amplifying neglect of other images and breaking cross-image reasoning chains [42].
To overcome this bottleneck, we propose IMAgent, the first open-source vision agent trained end-to-end with RL for multi-images environments. Architecturally, as shown in Figure 1, we expand the action space to support explicit visual indexing, enabling robust cross-image understanding. Cognitively, inspired by visual forgetting, we equip IMAgent with a collaborative visual function suite: visual confirmation (zoom_in for local details) and visual reflection (lookback_reuse to restore global context). Empirical analysis reveals that, without active tool use during chain-of-thought (CoT) reasoning, the model’s attention to input images gradually diminishes, especially in multi-image scenarios. As shown in Figure 2, token-level focus on image content drops to near zero in later inference stages. But IMAgent, through timely tool calls, refocuses image attention. Our pure RL framework allows the model to learn when, how, and which image index to interact with, establishing a robust “observe–confirm/reflect–iterate” paradigm.
Eliciting multi-turn tool use in multi-image settings is non-trivial. Single-image agents can often rely on supervised fine-tuning (SFT) to acquire reasoning patterns, but in an expanded action space, SFT leads to severe policy anchoring and high costs. To avoid blind exploration and to regularize the exploding multimodal action space for stable pure RL post-training, we introduce a Two-Level Mask strategy (action-level and trajectory-level), together with a correctness-coupled incentive mechanism.
Furthermore, existing work mainly focuses on using tools to improve model performance, either through SFT using carefully curated trajectory data [48], modifying RL algorithms [4, 74, 51], or combining these two methods [40, 39, 68]. Few studies explore the underlying principles by which tool use actually enhances agent performance. Based on the visual forgetting phenomenon in CoT, we conduct quantitative and qualitative analysis from an attention perspective, revealing for the first time that tool use can help agents focus on task-relevant regions in a timely manner, thereby improving performance.
To fill the data gap, we curate MIFG-QA, a 10k-sample fine-grained multi-image dataset for tool-interaction elicitation. Extensive experiments show IMAgent maintains competitive single-image performance while significantly improving multi-image fine-grained reasoning, and demonstrates encouraging generalization on other multi-image benchmarks. Beyond accuracy, analyses show that the reasoning behavior after training repeatedly bring the model back to task-relevant visual evidence, validating the underlying mechanism.
Figure 3: Overview of IMAgent. Our model will automatically choose whether and how to use tools based on the actual problem. Visual confirmation is achieved by cropping local regions of the image using zoom_in, and visual reflection is achieved by reusing the input image using lookbask_reuse.
@@
  • We propose IMAgent, the first multi-image, fine-grained, long-chain reasoning agent trained via end-to-end agentic RL, achieving dynamic cross-image coordination.
  • We identify systematic deficiencies in multi-image scenarios, specifically visual forgetting and the lack of explicit visual indexing. To address this, We design collaborative visual tools and a two-level mask strategy, and firstly reveal how tool use influences changes in visual attention.
  • We construct the MIFG-QA dataset to bridge the gap in multi-image fine-grained datasets. Extensive experiments on both single and multi-image datasets validate the state-of-the-art (SOTA) performance of IMAgent.
@@ 2.1 Multi-modal Reasoning Models
Multimodal reasoning models show a promising trend toward unifying multiple modalities within a language-centric framework[28]. Early works[55, 56, 58] focus on enhancing CoT reasoning through SFT and planning algorithms like Monte Carlo Tree Search[54], improving interpretability and reliability. Building on these foundations, RL-based methods shift the paradigm from imitating static reasoning patterns to actively optimizing them via feedback. Pioneering methods[60, 66, 9] apply Direct Preference Optimization[35] for behavioral alignment. After the success of Deepseek-R1[8], Group Relative Policy Optimization (GRPO)[37] becomes a mainstream algorithm, and Reinforcement Learning with Verifiable Rewards (RLVR) emerges as a powerful new training paradigm. These methods rapidly expand from mathematical and geometric reasoning[33, 63, 75] to visual skills[32, 31, 47], and further to modalities such as video[11] and audio[73]. Their success highlights the importance of high-quality reward signals, making data augmentation[53, 29], reward design[69, 49, 57, 65] and other modifications[62, 17] central to performance. Ultimately, RLVR drives breakthroughs in reasoning depth, coherence, and domain adaptability, laying a foundation for training advanced vision agents.
@@ 2.2 Vision Agents
Vision Agents aim to break the limitations of static, parameterized knowledge by autonomously using tools, deeply integrating dynamic visual perception with action-based reasoning[5]. Early approaches relied on prompt engineering to passively trigger tools[21, 43, 14, 61], while current mainstream methods increasingly adopt RL post-training, driving advances in multimodal retrieval[50, 15, 45] and integrated code execution[16, 44, 26, 10]. Notably, OpenAI’s O3 and Qwen3-VL[41] enable models to “think with images”, generating intertwined chains of thought that combine text and visuals. Inspired by O3’s great performance, many open-source solutions have emerged[74, 4, 68, 24, 39, 67], typically adopting a two-stage post-training process: SFT for tool call format learning, followed by RL to enhance visual tool use for complex tasks. However, this approach requires carefully curated SFT datasets and limits generalization. Deepeyes[74] cultivates tool-calling ability through pure reinforcement learning, but is limited to single image input and a single tool type. Though, aforementioned works achieve partial power of O3, they mainly focus on reproducing the tool use ability in single-image scenarios, largely neglecting the more common multi-image input settings. Addressing this gap is one of the main motivations of our work.
Figure 4: Action-Level and Trajectory-Level Mask.
@@ 3.1 Reasoning Pattern
Adhering to the “observe–confirm/reflect–iterate” paradigm, IMAgent outputs the final answer after obtaining sufficient information. Specifically, the complete reasoning process can be described as follows:
@@
    Initial Observation. Given an image set $I$ and a corresponding question $Q$, IMAgent performs an initial reasoning step. During this stage, the model autonomously decides whether to invoke visual confirmation or reflection. If the model calls a tool, it will output structured tool use tags.
@@
    Calling Visual Tools. When the model determines that there is a need for a visual tool call, it will generate the tool name and the required parameters. For visual confirmation via local image cropping, the model outputs an image index to specify the target image, a bounding box to indicate the region to zoom in on, and a label to emphasize the region of interest. For visual reflection via image reuse, the model outputs an image index and a reflection reason, which reinforces the purpose of reflection and enhances logical interpretability.
@@
    Observation and further reasoning. The returned results are concatenated with the historical context in an appropriate format, forming the input for the current step. The model continues reasoning based on this augmented trajectory.
@@ 3.2.2 Two-Level Mask Strategy
Action-Level Mask.
After the model issues a tool call request, the execution results of the tool (including the execution status and returned images) are concatenated into the previous context and used as input for the next round of reasoning. However, since this information often deviates from the distribution of existing knowledge, it can lead to abnormal directions in parameter optimization. To mitigate this issue, we apply a mask to information other than the model’s output to prevent it from participating in policy updates.
Trajectory-Level Mask. During training, we observe that some interaction trajectories produce excessively long outputs or exceed the predefined maximum interaction rounds. These issues induce training instability and impede the model from learning effective tool-usage strategies. To address this problem, we propose a trajectory-level masking mechanism to filter out such invalid trajectories.
Specifically, we evaluate the $N$ trajectories generated for each sample in the GRPO algorithm. If a trajectory is identified as invalid, the entire trajectory is masked and excluded from policy updates. This mechanism introduces nearly no additional computational overhead, since it directly masks existing generated trajectories rather than generating new valid ones. In this work, a trajectory is marked invalid if it exceeds the maximum interaction iterations, surpasses the maximum response length (typically caused by repetitive model outputs), or fails to produce any final answer. This mask also avoids reward-induced suppression of the model’s long-range exploration.
3.3 Curation of MIFG-QA
To train the model’s dynamic visual reasoning capabilities for multi-image tasks, we design a three-stage data synthesis pipeline (as shown in Figure 5). Stage 1: Images Selection. Considering the necessity of activating visual capabilities, we exclude low-resolution images and selected data with rich visual content. Specifically, for natural images, we randomly sample high-resolution data from MGrounding [27], covering image groups with relationships such as difference, contrast, and time.
For scientific figures, we constructe the dataset from high-resolution conference posters [36]. We perform adaptive structural segmentation on individual conference posters, treating different segmented regions as multi-image data. The resulting multi-image dataset maintains strong contextual relevance.
@@ 4.1 Experimental Setup
Table 1: Results of different models on benchmarks. Bold text indicates the best score.
Model No SFT Single-Image Benchmarks Multi-Image Benchmarks
V* HR-4K HR-8K MME Avg MIFG-QA BLINK Mantis M4Bench Avg
GPT-4o [19] 66 65 59.6 62.8 63.4 48.11 60
Gemini2.5-Flash [7] 72.3 77.5 73.7 60.9 71.1
Gemini2.5-pro [7] 79.1 83.9 81.5 71.3 79.0
Qwen3-VL-235B [2] 80.6 83.0 80.4 71.6 78.9
Qwen2.5-VL-7B [3] 77.0 70.6 66.7 57.3 67.9 41.98 55.5 75.6 45.2 54.6
[…]
SEAL [52] 75.4
Dyfo [25] 81.2
Deepeyes [74] 85.6 75.0 69.8 64.1 73.6 38.36 55.7 75.1 54.6 56.1
Pixel-Reasoner [39] 84.3 72.6 66.1 64.4 71.9 44.49 56.3 78.1 42.8 55.6
Deepeyes-V2 [18] 81.8 77.9 73.8 64.9 74.6
Thyme [70] 82.2 77.0 72.0 64.8 74.0 56.1
SenseNova-Mars-7B [6] 79.1 69.1
IMAgent-7B 88.5 74.0 71.5 64.9 74.7 49.21 56.7 78.8 56.5 60.0
Qwen3-VL-8B [2] 86.4 78.9 74.6 61.9 75.5 47.17 66.7 77.8 47.4 59.8
Qwen3-VL-30B [2] 82.2 78.5 74.2 67.7 75.7 48.11 66.4 77.8 53.4 61.4
Skywork-R1V4(30B) [72] 88.0 82.8 79.8 71.4 80.3
SenseNova-Mars-8B [6] 92.2 83.1 78.4 67.9 80.4 50.47 66.3 77.8 54.2 62.2
IMAgent-8B 90.6 85.3 82.9 70.5 82.3 53.86 67.8 78.3 54.7 63.7
@@
Hyperparameters. We adopt an end2end reinforcement learning algorithm, GRPO, with Qwen3-VL-8B-Instruct as the base model. The average number of pixels per sample in our MIFG-QA (poster) exceeds 9 million, which occupies over 12k context tokens. We therefore set the maximum pixel budget to 4 million. This preserves context space to accommodate more interaction turns. The model is trained on H20 GPUs with a batch size of 256, 8 rollouts per prompt, and a maximum limit of 5 tool interactions. The KL coefficient is set to 0.0, with the maximum input length restricted to 10,480 tokens and the maximum response length to 20,480 tokens. Main experimental results are presented in Table 1, where the values are the average of repeated experiments, and the maximum pixel limit of input images is 16384×28×28.
@@
Baselines and Benchmarks. We conduct a systematic comparison of IMAgent against advanced proprietary models (e.g., GPT-4o [19]), open-source models (e.g., Qwen2.5-VL [3]), and models implementing multimodal CoT (e.g., Deepeyes [74], Pixels Reasoner [39], SenseNova [6]).
@@
For evaluation benchmarks, we perform comprehensive comparisons on three classic single-image fine-grained benchmarks (V* Bench [52], HR-Bench [46], and MME-RealWorld [71]) and four multi-image datasets (BLINK [12], Mantis [20], M4Bench [59], and MIFG-QA-val) to validate the effectiveness of IMAgent’s dynamic visual reasoning. The test data of V* Bench and HR-Bench includes ultra-high-resolution images with resolutions ranging from 2k to 8k, where the target objects relevant to the test questions account for a very small proportion of pixels. MME-RealWorld is a large-scale high-resolution dataset of real-world scenarios, covering 43 scenarios, which focuses on evaluating the perception and reasoning capabilities of multimodal large models. BLINK [12] and Mantis [20] are general multi-image benchmarks, while M4Bench [59] and MIFG-QA-val are benchmarks that focus on evaluating fine-grained understanding.